# Deep Unfolding-Aided Parameter Tuning for Plug-and-Play Based Video Snapshot Compressive Imaging

### v2

@@ V-C Evaluation for Real Data
Finally, we evaluate the performance of the proposed parameter tuning for real data [23]. Figures 8 and 9 show the reconstruction results of PnP-ADMM for a real video domino and water balloon, respectively.
@@
In the figures, the top row shows the reconstruction results for each frame of conventional first parameter setting in Section III-C, and the bottom row shows the reconstruction results of the proposed parameters. In the experiments, $N_{t}=10$ frames of the video are reconstructed from a single compressed image. To obtain better reconstruction results, we use the initialization with GAP-TV for the real data. From Figs. 8 and 9, we can see that the proposed parameter tuning can achieve better reconstruction results than the conventional ones for real data, even though the parameter is trained by using the synthetic data. However, since there are still some artifacts in the reconstruction results (e.g., the latter frames in the domino video), further improvement would be required for the real data. One possible approach is to use the joint denoising strategy [23], where the weighted sum of multiple denoiser is used in the denoinsing process. Although we do not focus on the joint denoising strategy in this paper, the proposed parameter tuning can also be combined with the strategy.
@@ VI Conclusion
In this study, we have focused on the design of the noise level paremters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ in the PnP-based video SCI algorithms. We have proposed the data-driven tuning on the basis of the deep unfolding approach. In the proposed method, we prepare the dataset for the video SCI and train the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ by minimizing the loss function. The simulation results show that the proposed noise level parameters can achieve higher reconstruction accuracy than the conventional ones. Moreover, the trained parameters can achieve good performance even when the standard deviation of the measurement noise is different from that in the training data. Since the performance improvement is obtained by the choice of the noise level parameters only, our results demonstrate the importance of the parameter tuning in PnP-based video SCI.
Future work includes the training of other parameters like $\rho$ in PnP-ADMM, examination of the robustness to the mask, and extension for the reconstruction of color videos. The integration of the proposed parameter tuning with other SCI reconstruction methods is also an interesting topic for further development of SCI.

### v3

@@ V-C Evaluation for Real Data
Finally, we evaluate the performance of the proposed parameter tuning method for real data [23]. Figures 10 and 11 show the reconstruction results of PnP-ADMM for a real video domino and water balloon, respectively.
@@
In the figures, the top three rows show the reconstruction results for each frame of the conventional first parameter setting (step) in Section III-C, and the bottom row shows the reconstruction results of the proposed parameters. For the conventional ones, we show the results for $\rho=0.0001,0.01$, and $10$. For the proposed method, we show only the results with $\rho=0.01$ because the training is performed with $\rho=0.01$ as described in Section IV-B3. In the experiments, $N_{t}=10$ frames of the video are reconstructed from a single compressed image. To obtain better reconstruction results, we use the initialization with GAP-TV for the real data. From Figs. 10 and 11, we can see that the proposed parameter tuning can achieve better reconstruction results than the conventional ones for real data, even though the parameter is trained by using the synthetic data. However, since there are still some artifacts in the reconstruction results (e.g., the latter frames in the domino video), further improvement would be required for the real data. One possible approach is to use the joint denoising strategy [23], where the weighted sum of multiple denoisers is used in the denoinsing process. Although we do not focus on the joint denoising strategy in this paper, the proposed parameter tuning can also be combined with the strategy.
@@ VI Conclusion
In this study, we have addressed the critical challenge of designing noise level parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ in PnP-based video SCI algorithms. To this end, we have proposed a data-driven tuning approach leveraging the deep unfolding framework, which systematically learns the parameters by minimizing a loss function derived from the reconstruction task. Through extensive simulations, we demonstrated that the proposed method outperforms the conventional manually designed parameters in the reconstruction accuracy. Remarkably, the trained parameters maintained robust performance even under varying levels of measurement noise, highlighting their adaptability and effectiveness. Since the performance improvement is obtained by the choice of the noise level parameters only, our results demonstrate the importance of the parameter tuning in PnP-based video SCI.
An intriguing finding of this study is the non-monotonic behavior of the learned noise level parameters, which diverges from the conventional assumption of monotonicity required for several convergence analyses in PnP-based methods. This result implies that higher reconstruction accuracy can be achieved with parameter settings that deviate from these assumptions. To bridge this gap, future work should focus on developing theoretical frameworks to explain the effectiveness of non-monotonic parameter sequences and exploring adaptive tuning strategies that dynamically optimize parameters during the iterative process. Such efforts could enhance the robustness and applicability of the PnP methods.
Future work also includes the training of other parameters such as $\rho$ in PnP-ADMM, examination of the robustness to the mask, and extension for the reconstruction of color videos. The integration of the proposed parameter tuning with other SCI reconstruction methods is also an interesting topic for further development of SCI. Addressing these challenges would improve the generality and applicability of the proposed method in practical scenarios.